OpenStudy (anonymous):

A runner starts from rest and speeds up with constant acceleration. If she has gone a distance of 30 m at the point when she reaches a speed of 8 m/s, what is her acceleration?

OpenStudy (akashdeepdeb):

s = 30m u = 0m/s a = ? v = 8m/s Use the formula: \[a = \frac{v^2-u^2}{2s}\] Got it? :)

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!